OpenStudy (anonymous):

What value could be added to 2/15 to make the sum greater than 1/2

OpenStudy (anonymous):

at least 6/15

OpenStudy (anonymous):

d 8/15

OpenStudy (anonymous):

\[x+\frac{2}{15}>\frac{1}{2}\] \[x>\frac{1}{2}-\frac{2}{15}\] \[x>\frac{11}{30}\]